OpenStudy (swissgirl):

Define \( \oplus \text{ on } \mathbb{R} X \mathbb{R} \) by setting \( (a,b) \oplus (c,d)=(ac-bd, ad+bc) \) Show that the function h from system \( (\mathbb{C}, \cdot) \to\ (\mathbb{R} x \mathbb{R}, \oplus) \) given by h(a+bi)=(a,b) is a one to one function from the set of complex numbers that is onto \(\mathbb{R} x \mathbb{R}\) and is operation perserving

OpenStudy (swissgirl):

I have no background in complex numbers whatsoever.

OpenStudy (anonymous):

then this is going to be a treat

OpenStudy (anonymous):

but as i wrote before this is how you construct the complex numbers out of the real numbers the map you want is a simple one, takes \(a+bi\to (a,b)\)

OpenStudy (swissgirl):

right u started explaining it b4 but i so didnt follow cuz i never really studied it

OpenStudy (anonymous):

it is certainly onto , and there is not much to show here what you have to do is pick an element of \(\mathbb{R}\times \mathbb{R}\) and show that it comes from some complex number, but it is so simple as to almost be confusing if \((a,b)\in \mathbb{R}\times \mathbb{R}\) then it comes from the complex number \(a+bi\) that is \(h(a+bi)=(a,b)\) and so it is onto

OpenStudy (swissgirl):

so every complex number comes from a real number meaning \( \mathbb{C} \subseteq \mathbb{R} \) ?

OpenStudy (anonymous):

oh no this is not \(\mathbb{R}\)

OpenStudy (anonymous):

it is \(\mathbb{R}\times \mathbb{R}\)

OpenStudy (swissgirl):

ohhh right

OpenStudy (anonymous):

it is just a way of constructing them

OpenStudy (anonymous):

one to one is easy too, because \(a+bi=c+di\iff a=c, b=d\)

OpenStudy (anonymous):

that is, two complex numbers are equal iff their real parts and imaginary parts are equal

OpenStudy (swissgirl):

right got that

OpenStudy (anonymous):

so if \(h(a+bi)=h(c+di)\) that means \((a,b)=(c,d)\) which in turn means \(a=c\) and \(b=d\) and so \(a+bi=c+di\)

OpenStudy (anonymous):

last thing you have to show is \[h((a+bi)(c+di))=h(a+bi)\oplus(c+di)\]

OpenStudy (anonymous):

where the multiplication on the left is complex number multiplication, and the multiplication on the right is the \(\oplus\) multiplication defined earlier it work for sure

OpenStudy (swissgirl):

Thanks @satellite73

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

you do know how to multiply two complex numbers together right? that is all that is really needed here

OpenStudy (swissgirl):

hahah nooo but ill research that

OpenStudy (anonymous):

really?

OpenStudy (swissgirl):

I have never touched complex numbers in my life

OpenStudy (swissgirl):

idk in HS we never covered it and this is the first time in university that it is coming up

OpenStudy (anonymous):

since \(i=\sqrt{-1}\) you get \[(a+bi)(c+di)=(ac-bd)+(ac+bd)i\] pretty much identical to what you have on the right, with the plus sign in the middle replaced by a comma

OpenStudy (anonymous):

ok that was a typo you get \[(a+bi)(c+di)=(ac-bd)+(ad+bc)i\]

OpenStudy (swissgirl):

(a+bi)(c+di)=(ac−bd)+(ad+bc)i y is the i only after the second term?

OpenStudy (swissgirl):

ohhhh i get it u just grouped it like that

OpenStudy (swissgirl):

Thanks satellite for having patience with me

OpenStudy (swissgirl):

didnt go to a good hs so i am missing the basics in math whtvr

OpenStudy (anonymous):

you ca multiply out using the distributive law as well and you get the term \(bdi^2\) but since \(i^2=-1\) it sure in to \(-bd\)

OpenStudy (swissgirl):

gotcha Thanks :DDDDDDDDDDDDDDDDD

OpenStudy (anonymous):

yw